OpenStudy (anonymous):

what is the remainder of 7x^2+5x−24 divided by (x−2). A. 0 B. -8 C. 14 D. 3

OpenStudy (raden):

first, solve for x if x-2=0 get x = 2 now subtitute x = 2 into 7x^2+5x−2. it means 7(2)^2+5(2)−2 = 7(4) + 10 - 2 = 28 + 10 - 2 = 36